# Exploring the quantum capacity of a Gaussian random displacement channel using Gottesman-Kitaev-Preskill codes and maximum likelihood decoding

### v1

@@ Abstract
Approximate MLD for color-hexagonal GKP codes using a tensor-network decoder–As shown in Fig. 1, the color-hexagonal GKP code family sometimes outperforms the surface-square GKP code family in terms of the achievable rate. We thus provide additional numerical data on the color-hexagonal GKP codes. In Fig. 3, we show the fidelities of the color-hexagonal GKP code using an approximate MLD and the surface-square GKP code using the exact MLD. The approximate MLD for the color-hexagonal GKP code is implemented by using a tensor-network decoder Chubb (2021); Chubb and Lapeyre (2022) with a maximum bond dimension of $64$ during tensor-network contractions. We observe that despite the suboptimal, approximate MLD, the color-hexagonal GKP code family outperforms the surface-square GKP code family in the regime of $0.602\leq\sigma\leq 0.607$ at the same respective code distance $d$. When combined with the promising asymptotic trend of the surface-square GKP code family shown in Fig. 2(b), the numerical results in Fig. 3 suggest that the asymptotic performance (in the $d\rightarrow\infty$ limit) of the color-hexagonal GKP code may have an important implication on the quantum capacity of the Gaussian random displacement channel. It is thus an interesting open question as to whether the error threshold of the color-hexagonal GKP code family can exceed $\sigma=1/\sqrt{e}$ or not (especially when an exact MLD is used) and correspondingly whether a non-zero quantum state transmission rate can be acheived for a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ with $\sigma\geq 1/\sqrt{e}$. A positive result in this direction will establish an improved lower bound of $Q(\mathcal{N}_{\sigma})$.
@@
Conclusion and outlook–In this work, we have considered an open problem of determining the quantum capacity of a Gaussian random displacement channel (i.e., $Q(\mathcal{N}_{\sigma})$). In particular we have aimed to make progress on this problem by analyzing concrete instances of multi-mode GKP codes using either an exact or approximate maximum likelihood decoder. Our results clearly illustrate the particular relevance of GKP codes which encode just one logical qubit over multiple bosonic modes when studying a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ in a high noise regime (e.g., $\sigma\sim 1/\sqrt{e}$).
@@
Note that the maximum likelihood decoding of a quantum error-correcting code can be often be mapped to a statistical mechanical model Dennis et al. (2002); Chubb and Flammia (2021) and this connection has been explored in the context of toric-GKP codes Vuillot et al. (2019) as well. An interesting avenue for future research is to extend this relationship further and attempt to connect this mapping also to the problem of quantum communication theory since our work reveals a close connection between the maximum likelihood decoding of the surface-square and color-hexagonal GKP codes and the problem of Gaussian quantum channel capacity.
Lastly, an earlier work Lin et al. (2023) provided several instances of numerically-optimized multi-mode GKP codes that outperform the known concatenated GKP codes, even including the surface-square GKP code and the color-hexagonal GKP code, with the same number of modes. The performances of these codes have only been analyzed through a suboptimal closest-point decoder. Thus, continued search of new multi-mode GKP codes beyond concatenation and their performance analysis using maximum-likelihood decoder can also be fruitful future research directions, potentially having relevance to the problem of Gaussian quantum channel capacity as well.
Acknowledgements–We would like to acknowledge the AWS EC2 resources which were used for part of the numerical simulations performed in this work. ML would like to thank Péter Kómár and Eric Kessler for their supports of the project. We also would like to thank the Day 1 Science Mentorship (D1SM) program led by Connor Stewart at Amazon through which this project was initiated.
Supplemental Material
@@ Appendix B Maximum likelihood decoding for GKP codes
In this section, we review MLD for GKP codes following Ref. Conrad et al. (2022). Let $\Lambda$ denote the lattice corresponding to the GKP code considered, where the lattice points $\boldsymbol{u}\in\sqrt{2\pi}\Lambda$ correspond to the stabilizers of the GKP code. The prefactor of $\sqrt{2\pi}$ ensures the stabilizers commute with each other. For the purpose of decoding, we will need to consider the symplectic dual lattice $\Lambda^{\perp}$ where the quotient
@@
By measuring the stabilizers, we obtain $\boldsymbol{s}\in\mathbb{R}^{2N}$, a set of real-valued syndromes. Since the true error $\boldsymbol{\xi}$ is not known a priori, we assume $\boldsymbol{\eta}_{\boldsymbol{s}}$ to be a candidate error that is consistent with the syndrome $\boldsymbol{s}$ (for more details on the definitions of syndrome and candidate errors, we refer the readers to Sec. IV A of Ref. Lin et al. (2023)). With these, MLD aims to find the logical operator $\boldsymbol{l}^{\perp}$ that maximizes the following coset probability Conrad et al. (2022)
@@
After the optimal solution $\boldsymbol{l}_{*}^{\perp}$ is found, the corresponding displacement operator will be applied to attempt to correct the random shift error $\boldsymbol{\xi}$. Note that the MLD is an optimal strategy to minimize the chance of getting a logical error after the attempted shift correction Bravyi et al. (2014).
@@ Appendix C More details on the exact MLD for the surface-square GKP codes
In this section, we review the exact MLD for the “unrotated” surface code, proposed by Bravy, Suchara and Vargo (BSV) in Ref. Bravyi et al. (2014). We will explain how to adapt this method to perform an efficient and exact MLD on a “rotated” surface-square GKP code.
In contrast to the rotated surface-square GKP code considered in our work, Ref. Bravyi et al. (2014) considered the surface code in a different geometry, which is often referred to as the “unrotated” surface code. In Fig. 6(a), we show an example of $d=5$ “unrotated” surface code where the black dots represent the data qubits in the code. In the “unrotated” surface code, the $X$-type stabilizers (red) are often referred to as the site stabilizers, which act on four data qubits of the vertices, and the $Z$-type stabilizers (green) are the plaquette stabilizers, which act on four data qubits of the square plaquettes. Stabilizers of either type that are on the boundaries act on only three data qubits. In our work, to study “rotated” surface-square GKP codes, we embed a “rotated” surface-square GKP code into an “unrotated” surface code as shown in Fig. 6(a). This is the starting point of applying the BSV decoder to surface-square GKP codes.
@@
One of the seminal contributions in Ref. Bravyi et al. (2014) is a method for calculating the coset probabilities exactly and efficiently for the “unrotated” surface codes subject to a pure $Z$-noise or $X$-noise model. More specifically, there are $N^{\prime}\equiv d^{2}+(d-1)^{2}$ data qubits for the distance-$d$ “unrotated” surface code and we assign to each data qubit a weight $\tau^{\prime}_{i}\in\mathbb{R}$ for $1\leq i\leq N^{\prime}$. Then, let $\boldsymbol{g}^{\prime}$ denote a binary vector with $N^{\prime}$ components describing an $X$-type stabilizer of the “unrotated” surface code. The coset probabilities for the “unrotated” surface code can then be written in the following form for a given set of weights $\boldsymbol{\tau}^{\prime}$,
@@
where $\mathcal{G}_{X}^{\prime}$ is the set of all the $X$-type stabilizers of the “unrotated” surface code, and $g^{\prime}_{i}$ denotes the $i$-th component of $\boldsymbol{g}^{\prime}$. Note that $(\tau^{\prime}_{i})^{g^{\prime}_{i}}=\tau^{\prime}_{i}$ if $g^{\prime}_{i}=1$ and $(\tau^{\prime}_{i})^{g^{\prime}_{i}}=1$ if $g^{\prime}_{i}=0$. Notably, $Z(\boldsymbol{\tau}^{\prime})$ can be calculated efficiently in time $O(d^{4})$ (see Sec. V of Ref. Bravyi et al. (2014) for more details). Upon adequate adaptations, this technique can be used to efficiently decode a surface-square GKP code as we describe below.
@@
As shown in Ref. Lin et al. (2023) (see Sec. IX therein), $\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$ contains $2\sqrt{\pi}Z_{N}$ as a sublattce, which allows us to classify the lattice points in $\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$ to different equivalent classes. In particular, for any $\boldsymbol{u}\in\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$, it can be written as $\boldsymbol{u}=\sqrt{\pi}(\boldsymbol{g}+\boldsymbol{v})$ where $\boldsymbol{g}$ is the binary vector for a $X$-type stabilizer, and $\boldsymbol{v}\in 2Z_{N}$. Then defining
@@
Here $\mathcal{G}_{X}$ denotes the group of $X$-type stabilizers. Equations similar to Eqs. (10)-(11) for general concatenated-square GKP codes have been obtained in Ref. Conrad et al. (2022), but the authors did not investigate further the performance of the MLD for explicitly GKP codes. Below we will explain how to use the technique in Ref. Bravyi et al. (2014) to calculate the quantity in Eqs. (10)-(11) for the specific case of the surface-square code, given its similarity to $Z(\boldsymbol{\tau}^{\prime})$ in Eq. (7).
@@ Appendix D More details on the tensor-network decoder for the color-GKP codes
We remark that Eqs. (18)-(19) is applicable for generic concatenated-GKP code. Because $S=I_{2}$ for the concatenated-square GKP code, Eq. (19) reduces to a product of two summations for the $\hat{q}$ and $\hat{p}$ subspaces, where each summation has the same form as Eq. (16). In other word, Eqs. (18)-(19) represent a straightforward generalization for the result obtained in Ref. Conrad et al. (2022), for approximating MLD for generic concatenated-GKP codes.
The tensor-network method to calculate the quantities in Eq. (16) and Eq. (18) was first described in Ref. Bravyi et al. (2014) for decoding surface code, and later generalized to general 2D stabilizer codes in Ref. Chubb (2021). The tensor network for the color-square code is depicted in Fig. 7(a) which typically involves two types of tensors. The first type of tensors, shown in Fig. 7(b), correspond to stabilizers and take value one iff all the indices of the tensor have the same value, and zero otherwise. The second type of tensors, shown in Fig. 7(c), correspond to qubits which, as defined in Eq. (11), depend on the values of the candidate error $\boldsymbol{\eta}_{\boldsymbol{s}}$, the logical error $\boldsymbol{l}^{\perp}$ and the stabilizer $\boldsymbol{g}$ at the qubit. Taking the rank-3 tensor in Fig. 7(c) as an example, we have $g_{i}=r\oplus s\oplus t$ where $r,s,t$ are the indices of the tensor and $g_{i},r,s,t\in\mathbb{F}_{2}\equiv\left\{0,1\right\}$. For the rest of this section, we use $\oplus$ to denote the addition in $\mathbb{F}_{2}$ or $\mathbb{F}_{4}$, see below. We note that $g_{i}$ is essentially the local logical error for the $i$-th qubit, if we view $\boldsymbol{\eta}^{\prime}$, defined in Eq. (9), as an effective candidate error.
@@ Appendix E More details on the brute force MLD for the $[[5,1,3]]$-hexagonal GKP code
In this section, we provide more details on the brute-force MLD for the $[[5,1,3]]$-hexagonal GKP code. The brute-force MLD aims to evaluate exactly the same expressions in Eq. (18)-19, except that $\mathcal{G}\equiv\mathcal{G}_{513}$ is for the full stabilizer group of the $[[5,1,3]]$ code. We have listed the generators for $\mathcal{G}_{513}$ in Tab. 1. In total there are $16$ elements in $\mathcal{G}_{513}$ which makes the brute-force MLD feasible. As an example, let us consider the summand for the stabilizer $IYXXY$ which is a product of five terms, each has the same form as that in Eq. (19). Taking second qubit as an example, because the second Pauli operator of the stabilizer is $Y$, the corresponding $[\boldsymbol{g}]_{i=2}=[1,1]^{T}$ using our convention. With that, we can plug each $[\boldsymbol{g}]_{i}$ into Eq. (19) to estimate the weight in the $i$-th qubit. Specifically, let
@@ Appendix F More discussions on prior works
In this section, we compare our results to several prior works Fukui et al. (2017, 2018); Hänggli et al. (2020); Zhang et al. (2023), which also study the thresholds of multi-mode GKP codes near $\sigma=1/\sqrt{e}\simeq 0.6065$.
@@
In Ref. Fukui et al. (2017), the authors considered the $C_{4}/C_{6}$ code concatenated with square-GKP codes. They showed that the code has threshold between $0.60$ and $0.61$ when the analog information from the inner square GKP codes is incorporated into the decoder. The numerical evidence is shown in Fig. 3 therein, where the noise strength $\sigma$ is scanned with the resolution of $0.01$. Because of the limited resolution, it is difficult to conclude that the code has a threshold above $\sigma=1/\sqrt{e}\simeq 0.6065$.
@@
In Ref. Fukui et al. (2018), the authors applied the decoding strategy in Ref. Fukui et al. (2017) to surface-square GKP code, and the results are presented in Fig. 2 therein. Close inspection of Fig. 2(b) reveals that surface-square GKP code at $d=25$ has higher failure probability than $d=23$ at $\sigma=0.605$ using the suboptimal decoding method therein. In order to resolve the ambiguities in Ref. Fukui et al. (2017, 2018), we have chosen to scan the noise strength with higher resolution of 0.001 in our simulations. Moreover, we have carefully studied the fidelity of the surface-square GKP code near $\sigma=1/\sqrt{e}\approx 0.6065$ as shown in Fig. 2(b) in the main text. The result in Ref. Fukui et al. (2018) is consistent with our conclusion that surface-square code can have threshold near $\sigma=1/\sqrt{e}\approx 0.6065$ but only by using the optimal decoding method.
@@
In Ref. Hänggli et al. (2020), the authors studied the QEC capacity of several multi-mode GKP codes, including surface-rectangular and surface-asymmetric-hexagonal GKP codes. With a tensor-network decoder, they also showed that the analog information from the inner GKP codes can significantly improve the QEC capacity of these codes, and the main numerical results are shown in Fig. 8 therein. But, similar to Ref. Fukui et al. (2017), the noise strength $\sigma$ is scanned only with the resolution of $0.01$, and one can only conclude that these codes have threshold between $0.60$ and $0.61$. We also note that the decoders in Ref. Hänggli et al. (2020) are approximate MLD since a tensor-network decoder with a finite maximum bond dimension cutoff was used. However as we showed above, MLD of a surface-rectangular GKP code can in fact be performed exactly and efficiently with our adaptation of the BSV method because the shift errors in the position and momentum subspaces are decoupled from each other. Hence, the threshold of the surface-rectangular GKP code shown in Fig. 8(b) in Ref. Hänggli et al. (2020) is not necessarily optimal. It would be interesting to use our adaptation of the exact BSV method to explore the thresholds of the the surface-rectangular GKP codes with various aspect ratios, or surface-GKP codes on rectangular geometries (instead of square geometry shown in Fig. 6).
@@
In Ref. Zhang et al. (2023), the authors studied the XZZX-rectangular GKP codes decoded by using a minimum-weight-perfect-matching decoder. In Fig. 4 therein, the authors claimed that the code has threshold of $0.67$ when the square root of the aspect ratio of the inner rectangular GKP codes is equal to $2.1$. We note, however, that the simulation was only carried out for codes up to distance $d=9$. Further, close inspection of Fig. 4(b) reveals that the crossings of the logical error rates for distances $d$ and $d+2$ decreases as the distance is increased from $d=3$ to $d=7$. It is reasonable to suspect that the crossings could further decrease as the distance is increased further, and the threshold, a quantity defined only in the asymptotic limit of $d\rightarrow\infty$, could be much smaller than $0.67$ if the threshold ever exists for these codes. Hence, based only on the data presented, it is difficult to confidently conclude that the threshold of the XZZX-rectangular GKP code is larger than $\sigma=1/\sqrt{e}\simeq 0.6065$.

### v2

@@ V Approximate MLD for color-hexagonal GKP codes using a tensor-network decoder
As shown in Fig. 1, the color-hexagonal GKP code family sometimes outperforms the surface-square GKP code family in terms of the achievable rate. We thus provide additional numerical data on the color-hexagonal GKP codes. In Fig. 3, we show the fidelities of the color-hexagonal GKP code using an approximate MLD and the surface-square GKP code using the exact MLD. The approximate MLD for the color-hexagonal GKP code is implemented by using a tensor-network decoder [50, 52] with a maximum bond dimension of $64$ during tensor-network contractions. We observe that despite the suboptimal, approximate MLD, the color-hexagonal GKP code family outperforms the surface-square GKP code family in the regime of $0.602\leq\sigma\leq 0.607$ at the same respective code distance $d$. When combined with the promising asymptotic trend of the surface-square GKP code family shown in Fig. 2(b), the numerical results in Fig. 3 suggest that the asymptotic performance (in the $d\rightarrow\infty$ limit) of the color-hexagonal GKP code may have an important implication on the quantum capacity of the Gaussian random displacement channel. It is thus an interesting open question as to whether the error threshold of the color-hexagonal GKP code family can exceed $\sigma=1/\sqrt{e}$ or not (especially when an exact MLD is used) and correspondingly whether a non-zero quantum state transmission rate can be acheived for a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ with $\sigma\geq 1/\sqrt{e}$. A positive result in this direction will establish an improved lower bound of $Q(\mathcal{N}_{\sigma})$.
@@
VI Conclusion and outlook
In this work, we have considered an open problem of determining the quantum capacity of a Gaussian random displacement channel (i.e., $Q(\mathcal{N}_{\sigma})$). In particular we have aimed to make progress on this problem by analyzing concrete instances of multi-mode GKP codes using either an exact or approximate maximum likelihood decoder. Our results clearly illustrate the particular relevance of GKP codes which encode just one logical qubit over multiple bosonic modes when studying a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ in a high noise regime (e.g., $\sigma\sim 1/\sqrt{e}$).
@@ VI Conclusion and outlook
Note that the maximum likelihood decoding of a quantum error-correcting code can often be mapped to a statistical mechanical model [55, 56] and this connection has been explored in the context of toric-GKP codes [42] as well. An interesting avenue for future research is to extend this relationship further and attempt to connect this mapping also to the problem of quantum communication theory since our work reveals a close connection between the maximum likelihood decoding of the surface-square and color-hexagonal GKP codes and the problem of Gaussian quantum channel capacity.
Lastly, an earlier work [48] provided several instances of numerically-optimized multi-mode GKP codes that outperform the known concatenated GKP codes, even including the surface-square GKP code and the color-hexagonal GKP code, with the same number of modes. The performances of these codes have only been analyzed through a suboptimal closest-point decoder. Thus, continued search of new multi-mode GKP codes beyond concatenation and their performance analysis using maximum-likelihood decoder can also be fruitful future research directions, potentially having relevance to the problem of Gaussian quantum channel capacity as well.
VII Acknowledgements
We would like to acknowledge the AWS EC2 resources which were used for part of the numerical simulations performed in this work. ML would like to thank Péter Kómár and Eric Kessler for their supports of the project. We also would like to thank the Day 1 Science Mentorship (D1SM) program led by Connor Stewart at Amazon through which this project was initiated.
@@ Appendix B Maximum likelihood decoding for GKP codes
In this section, we review MLD for GKP codes following Ref. [46]. Let $\Lambda$ denote the lattice corresponding to the GKP code considered, where the lattice points $\boldsymbol{u}\in\sqrt{2\pi}\Lambda$ correspond to the stabilizers of the GKP code. The prefactor of $\sqrt{2\pi}$ ensures the stabilizers commute with each other. For the purpose of decoding, we will need to consider the symplectic dual lattice $\Lambda^{\perp}$ where the quotient
@@
By measuring the stabilizers, we obtain $\boldsymbol{s}\in\mathbb{R}^{2N}$, a set of real-valued syndromes. Since the true error $\boldsymbol{\xi}$ is not known a priori, we assume $\boldsymbol{\eta}_{\boldsymbol{s}}$ to be a candidate error that is consistent with the syndrome $\boldsymbol{s}$ (for more details on the definitions of syndrome and candidate errors, we refer the readers to Sec. IV A of Ref. [48]). With these, MLD aims to find the logical operator $\boldsymbol{l}^{\perp}$ that maximizes the following coset probability [46]
@@
After the optimal solution $\boldsymbol{l}_{*}^{\perp}$ is found, the corresponding displacement operator will be applied to attempt to correct the random shift error $\boldsymbol{\xi}$. Note that the MLD is an optimal strategy to minimize the chance of getting a logical error after the attempted shift correction [49].
@@ Appendix C More details on the exact MLD for the surface-square GKP codes
In this section, we review the exact MLD for the “unrotated” surface code, proposed by Bravy, Suchara and Vargo (BSV) in Ref. [49]. We will explain how to adapt this method to perform an efficient and exact MLD on a “rotated” surface-square GKP code.
In contrast to the rotated surface-square GKP code considered in our work, Ref. [49] considered the surface code in a different geometry, which is often referred to as the “unrotated” surface code. In Fig. 6(a), we show an example of $d=5$ “unrotated” surface code where the black dots represent the data qubits in the code. In the “unrotated” surface code, the $X$-type stabilizers (red) are often referred to as the site stabilizers, which act on four data qubits of the vertices, and the $Z$-type stabilizers (green) are the plaquette stabilizers, which act on four data qubits of the square plaquettes. Stabilizers of either type that are on the boundaries act on only three data qubits. In our work, to study “rotated” surface-square GKP codes, we embed a “rotated” surface-square GKP code into an “unrotated” surface code as shown in Fig. 6(a). This is the starting point of applying the BSV decoder to surface-square GKP codes.
@@
One of the seminal contributions in Ref. [49] is a method for calculating the coset probabilities exactly and efficiently for the “unrotated” surface codes subject to a pure $Z$-noise or $X$-noise model. More specifically, there are $N^{\prime}\equiv d^{2}+(d-1)^{2}$ data qubits for the distance-$d$ “unrotated” surface code and we assign to each data qubit a weight $\tau^{\prime}_{i}\in\mathbb{R}$ for $1\leq i\leq N^{\prime}$. Then, let $\boldsymbol{g}^{\prime}$ denote a binary vector with $N^{\prime}$ components describing an $X$-type stabilizer of the “unrotated” surface code. The coset probabilities for the “unrotated” surface code can then be written in the following form for a given set of weights $\boldsymbol{\tau}^{\prime}$,
@@
where $\mathcal{G}_{X}^{\prime}$ is the set of all the $X$-type stabilizers of the “unrotated” surface code, and $g^{\prime}_{i}$ denotes the $i$-th component of $\boldsymbol{g}^{\prime}$. Note that $(\tau^{\prime}_{i})^{g^{\prime}_{i}}=\tau^{\prime}_{i}$ if $g^{\prime}_{i}=1$ and $(\tau^{\prime}_{i})^{g^{\prime}_{i}}=1$ if $g^{\prime}_{i}=0$. Notably, $Z(\boldsymbol{\tau}^{\prime})$ can be calculated efficiently in time $O(d^{4})$ (see Sec. V of Ref. [49] for more details). Upon adequate adaptations, this technique can be used to efficiently decode a surface-square GKP code as we describe below.
@@
As shown in Ref. [48] (see Sec. IX therein), $\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$ contains $2\sqrt{\pi}Z_{N}$ as a sublattce, which allows us to classify the lattice points in $\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$ to different equivalent classes. In particular, for any $\boldsymbol{u}\in\sqrt{2\pi}\Lambda^{(\hat{p})}_{\text{surf-sq}}$, it can be written as $\boldsymbol{u}=\sqrt{\pi}(\boldsymbol{g}+\boldsymbol{v})$ where $\boldsymbol{g}$ is the binary vector for a $X$-type stabilizer, and $\boldsymbol{v}\in 2Z_{N}$. Then defining
@@
Here $\mathcal{G}_{X}$ denotes the group of $X$-type stabilizers. Equations similar to Eqs. (11)-(12) for general concatenated-square GKP codes have been obtained in Ref. [46], but the authors did not investigate further the performance of the MLD for explicitly GKP codes. Below we will explain how to use the technique in Ref. [49] to calculate the quantity in Eqs. (11)-(12) for the specific case of the surface-square code, given its similarity to $Z(\boldsymbol{\tau}^{\prime})$ in Eq. (8).
@@ Appendix D More details on the tensor-network decoder for the color-GKP codes
We remark that Eqs. (19)-(20) is applicable for generic concatenated-GKP code. Because $S=I_{2}$ for the concatenated-square GKP code, Eq. (20) reduces to a product of two summations for the $\hat{q}$ and $\hat{p}$ subspaces, where each summation has the same form as Eq. (17). In other word, Eqs. (19)-(20) represent a straightforward generalization for the result obtained in Ref. [46], for approximating MLD for generic concatenated-GKP codes.
The tensor-network method to calculate the quantities in Eq. (17) and Eq. (19) was first described in Ref. [49] for decoding surface code, and later generalized to general 2D stabilizer codes in Ref. [50]. The tensor network for the color-square code is depicted in Fig. 7(a) which typically involves two types of tensors. The first type of tensors, shown in Fig. 7(b), correspond to stabilizers and take value one iff all the indices of the tensor have the same value, and zero otherwise. The second type of tensors, shown in Fig. 7(c), correspond to qubits which, as defined in Eq. (12), depend on the values of the candidate error $\boldsymbol{\eta}_{\boldsymbol{s}}$, the logical error $\boldsymbol{l}^{\perp}$ and the stabilizer $\boldsymbol{g}$ at the qubit. Taking the rank-3 tensor in Fig. 7(c) as an example, we have $g_{i}=r\oplus s\oplus t$ where $r,s,t$ are the indices of the tensor and $g_{i},r,s,t\in\mathbb{F}_{2}\equiv\left\{0,1\right\}$. For the rest of this section, we use $\oplus$ to denote the addition in $\mathbb{F}_{2}$ or $\mathbb{F}_{4}$, see below. We note that $g_{i}$ is essentially the local logical error for the $i$-th qubit, if we view $\boldsymbol{\eta}^{\prime}$, defined in Eq. (10), as an effective candidate error.
@@ Appendix E More details on the brute force MLD for the $[[5,1,3]]$-hexagonal GKP code
In this section, we provide more details on the brute-force MLD for the $[[5,1,3]]$-hexagonal GKP code. The brute-force MLD aims to evaluate exactly the same expressions in Eq. (19)-(20), except that $\mathcal{G}\equiv\mathcal{G}_{513}$ is for the full stabilizer group of the $[[5,1,3]]$ code. We have listed the generators for $\mathcal{G}_{513}$ in Tab. 1. In total there are $16$ elements in $\mathcal{G}_{513}$ which makes the brute-force MLD feasible. As an example, let us consider the summand for the stabilizer $IYXXY$ which is a product of five terms, each has the same form as that in Eq. (20). Taking second qubit as an example, because the second Pauli operator of the stabilizer is $Y$, the corresponding $[\boldsymbol{g}]_{i=2}=[1,1]^{T}$ using our convention. With that, we can plug each $[\boldsymbol{g}]_{i}$ into Eq. (20) to estimate the weight in the $i$-th qubit. Specifically, let
@@ Appendix F More discussions on prior works
In this section, we compare our results to several prior works [39, 40, 44, 54], which also study the thresholds of multi-mode GKP codes near $\sigma=1/\sqrt{e}\simeq 0.6065$.
@@
In Ref. [39], the authors considered the $C_{4}/C_{6}$ code concatenated with square-GKP codes. They showed that the code has threshold between $0.60$ and $0.61$ when the analog information from the inner square GKP codes is incorporated into the decoder. The numerical evidence is shown in Fig. 3 therein, where the noise strength $\sigma$ is scanned with the resolution of $0.01$. Because of the limited resolution, it is difficult to conclude that the code has a threshold above $\sigma=1/\sqrt{e}\simeq 0.6065$.
@@
In Ref. [40], the authors applied the decoding strategy in Ref. [39] to surface-square GKP code, and the results are presented in Fig. 2 therein. Close inspection of Fig. 2(b) reveals that surface-square GKP code at $d=25$ has higher failure probability than $d=23$ at $\sigma=0.605$ using the suboptimal decoding method therein. In order to resolve the ambiguities in Ref. [39, 40], we have chosen to scan the noise strength with higher resolution of 0.001 in our simulations. Moreover, we have carefully studied the fidelity of the surface-square GKP code near $\sigma=1/\sqrt{e}\approx 0.6065$ as shown in Fig. 2(b) in the main text. The result in Ref. [40] is consistent with our conclusion that surface-square code can have threshold near $\sigma=1/\sqrt{e}\approx 0.6065$ but only by using the optimal decoding method.
@@
In Ref. [44], the authors studied the QEC capacity of several multi-mode GKP codes, including surface-rectangular and surface-asymmetric-hexagonal GKP codes. With a tensor-network decoder, they also showed that the analog information from the inner GKP codes can significantly improve the QEC capacity of these codes, and the main numerical results are shown in Fig. 8 therein. But, similar to Ref. [39], the noise strength $\sigma$ is scanned only with the resolution of $0.01$, and one can only conclude that these codes have threshold between $0.60$ and $0.61$. We also note that the decoders in Ref. [44] are approximate MLD since a tensor-network decoder with a finite maximum bond dimension cutoff was used. However as we showed above, MLD of a surface-rectangular GKP code can in fact be performed exactly and efficiently with our adaptation of the BSV method because the shift errors in the position and momentum subspaces are decoupled from each other. Hence, the threshold of the surface-rectangular GKP code shown in Fig. 8(b) in Ref. [44] is not necessarily optimal. It would be interesting to use our adaptation of the exact BSV method to explore the thresholds of the the surface-rectangular GKP codes with various aspect ratios, or surface-GKP codes on rectangular geometries (instead of square geometry shown in Fig. 6).
@@
In Ref. [54], the authors studied the XZZX-rectangular GKP codes decoded by using a minimum-weight-perfect-matching decoder. In Fig. 4 therein, the authors claimed that the code has threshold of $0.67$ when the square root of the aspect ratio of the inner rectangular GKP codes is equal to $2.1$. We note, however, that the simulation was only carried out for codes up to distance $d=9$. Further, close inspection of Fig. 4(b) reveals that the crossings of the logical error rates for distances $d$ and $d+2$ decreases as the distance is increased from $d=3$ to $d=7$. It is reasonable to suspect that the crossings could further decrease as the distance is increased further, and the threshold, a quantity defined only in the asymptotic limit of $d\rightarrow\infty$, could be much smaller than $0.67$ if the threshold ever exists for these codes. Hence, based only on the data presented, it is difficult to confidently conclude that the threshold of the XZZX-rectangular GKP code is larger than $\sigma=1/\sqrt{e}\simeq 0.6065$.